QUESTION # 2
The following information pertains to Glass Works Company for 2014.
Budgeted direct-labour cost: --------------------------------75,000 hours at $16 per hour
Actual direct-labour cost: ------------------------------------80,000 hours at $17.50 per hour.
Budgeted Manufacturing Overhead: ---------------------$ 997,500
Actual Manufacturing Overhead:
Depreciation ------------------------------------------------- $ 240,000
Property Taxes---------------------------------------------------12,000
Indirect Labour --------------------------------------------------82,000
Supervisory Salaries --------------------------------------------200,000
Utilities------------------------------------------------------------ 59,000
Insurance ----------------------------------------------------------30,000
Rental of Space-------------------------------------------------- 300,000
Indirect Material --------------------------------------------------- 79,000
Required
a) Compute the firm’s predetermined overhead rate, which is based on direct-labour
hours. (3 marks)
b) Calculate the overapplied or underapplied overhead for 2014. (4 marks)
c) Prepare a journal entry to close out the Manufacturing Overhead Account into Cost of
Goods Sold. (3 marks)
Answer #1

1) Overhead rate = estimated overhead/estimated direct labor = 997500/75000 = 13.30 per hour

2) Actual overhead = 1002000

Applied overhead = 80000*13.30 = 1064000

Over applied overhead = 1064000-1002000 = 62000

3) Journal entry

Date account and explanation Debit Credit
Manufacturing overhead 62000
Cost of goods sold 62000